What is Integration? Discuss types of Integration with Example.


1
Expert's answer
2021-11-15T16:36:01-0500

The integral of a function "\int_a^b{f(x)dx}" gives the area under the curve between the limits x=a and x=b.

Integrating refers to the summation of discreet data and is a reverse process of differentiation.

1) integration by parts

This type of integration is applied when integrating between a product of two functions.

Example:

"\int x\ sec^2x\ dx"

"=x\int Sec^2 x\ dx-\int (\int Sec^2x\ dx)(1dx)"

"=x\ tan \ x-\int tan\ x dx"

"=x\ tanx +ln|cos\ x|+C"


2) integration by partial fractions

This type of integration is applied when the function to be integrated are rational proper fraction in the form "\frac{P(u)}{Q(u)}"

Example:

"\int{\frac{2x+3}{(x-3)(x+1)}dx}"

"\frac{2x+3}{(x-3)(x+1)}=\frac{A}{x-3}+\frac{B}{x+1}"

"\implies2x+3=A(x+1)+B(x-3)"

A+B=2

A-3B=3

We solve above to get A="\frac{9}{4}" and B="-\frac{1}{4}"

Now "\int{\frac{2x+3}{(x-3)(x+1)}dx}=\int{\frac{\frac{9}{4}}{x-3}dx}+\int{\frac{-\frac{1}{4}}{x+1}dx}"

"=\frac{9}{4}ln|x-3|-\frac{1}{4}ln|x+1|+C"